### v1

# AniMaker: Automated Multi-Agent Animated Storytelling with MCTS-Driven Clip Generation

Abstract.
Despite rapid advancements in video generation models, generating coherent, long-form storytelling videos that span multiple scenes and characters remains challenging. Current methods often rigidly convert pre-generated keyframes into fixed-length clips, resulting in disjointed narratives and pacing issues. Furthermore, the inherent instability of video generation models means that even a single low-quality clip can significantly degrade the entire output animation’s logical coherence and visual continuity. To overcome these obstacles, we introduce AniMaker, a multi-agent framework enabling efficient multi-candidate clip generation and storytelling-aware clip selection, thus creating globally consistent and story-coherent animation solely from text input. The framework is structured around specialized agents, including the Director Agent for storyboard generation, the Photography Agent for video clip generation, the Reviewer Agent for evaluation, and the Post-Production Agent for editing and voiceover, collectively realizing multi-character, multi-scene animation. Central to AniMaker’s approach are two key technical components: MCTS-Gen in Photography Agent, an efficient Monte Carlo Tree Search (MCTS)-inspired strategy that intelligently navigates the candidate space to generate high-potential clips while optimizing resource usage; and AniEval in Reviewer Agent, the first framework specifically designed for multi-shot animation evaluation, which assesses critical aspects such as story-level consistency, action completion, and animation-specific features by considering each clip in the context of its preceding and succeeding clips. Experiments demonstrate that AniMaker achieves superior quality as measured by popular metrics including VBench and our proposed AniEval framework, while significantly improving the efficiency of multi-candidate generation, pushing AI-generated storytelling animation closer to production standards.
storytelling animation, multi agents, MCTS, storyboard generation
copyright: noneconference: ; ;
Figure 1. A visual example of AniMaker generating compelling storytelling animation from narrative text. Our framework maintains consistent character appearance across scenes while delivering high-quality action representation for complex sequences. AniMaker seamlessly integrates adaptive shot scheduling with smooth transitions, ensuring narrative coherence throughout the animation.
1. Introduction
The emergence of large-scale multimodal models (Yang et al., 2024; Kong et al., 2024; Liu et al., 2024c; Li et al., 2025) has ignited significant interest in storytelling videos. In response, a new wave of methods (Kim et al., 2024; Wu et al., 2024; Dalal et al., 2025; Guo et al., 2025) including MINT Video and TTT-Video has emerged, aiming to generate the entire video in a single pass. However, these methods still face significant challenges when generating long videos spanning multiple scenes and characters, particularly in maintaining visual continuity, narrative coherence, and non-repetitive content. In contrast, recent advancements in single-clip video generation, exemplified by models like Wan (Wan et al., 2025), Vidu(Bao et al., 2024), and Pika (Pika Labs, 2025), have pushed generated video clips closer to cinematic quality. Due to shorter clip duration, these models utilize textual and visual prompts directly and efficiently, improving motion quality, semantic fidelity, and consistency with reference images. Consequently, utilizing these models to generate a storytelling video with multiple clips appears highly promising.
Existing storytelling video frameworks (Lin et al., 2023; He et al., 2024; Li et al., 2024; Xu et al., 2025; Wu et al., 2025; Xie et al., 2024), which implement the composition of multiple video clips, generally follow a fixed pipeline: script → scene images → video clips → final video composition. While this modular format can effectively generate multi-shot long videos, it also introduces several limitations: Firstly, existing methods typically map scene images to fixed-length clips. It forms a rigid and fragmented video construction, leading to disjointed transitions and unnatural pacing, severely hindering the expressive continuity crucial for complex or extended actions. Secondly, due to the inherent instability of video generation models, a single flawed clip can noticeably degrade the overall video quality.
To mitigate fragmentation, an intuitive approach is generating continuous clips conditioned on prior frames, but this compounds error propagation and quality degradation. Drawing parallels with professional filmmaking, we identify that existing methods overlook Best-of-N Sampling—generating multiple clip candidates and selecting the best ones. However, implementing this faces two challenges: prohibitive computational costs and inadequate automated evaluation mechanisms. Generating and evaluating multiple candidate clips per shot is computationally intensive, often relying on expensive commercial APIs or prolonged GPU inference. Current evaluation metrics like VBench (Xing et al., 2024) only assess individual clips and their internal consistency, neglecting critical elements such as cross-clip coherence, sequential motion quality, and animation-specific qualities in storytelling animation.
To address these challenges, we introduce AniMaker, a multi-agent framework with MCTS-driven clip generation. This multi-agent framework mirrors professional production, including the Director Agent for storyboards, Photography Agent for clip generation, Reviewer Agent for evaluating the quality of clip candidates, and Post-Production Agent for editing the entire videos. These agents collaborate to enable automated multi-character, multi-scene storytelling without manual pre- or post-processing. The core Photography and Reviewer agents interact with each other under an MCTS-Gen scheme during clip generation. Specifically, based on Monte Carlo Tree Search (MCTS), MCTS-Gen offers an efficient strategy for navigating the vast candidate space of video generation. It strikes a balance between broad exploration and computational efficiency by intelligently allocating more generation opportunities to promising clips while encouraging the exploration of unexplored regions. For Reviewer Agent, we introduce a comprehensive evaluation framework AniEval, specifically designed for multi-shot storytelling animation. AniEval advances beyond metrics like VBench by implementing retrospective evaluation with cross-clip contextual references. It evaluates critical dimensions—story consistency, action completion, and animation-specific attributes—by analyzing each clip in the context of its preceding and succeeding clips. This contextual assessment serves as the quality evaluation mechanism for video-clip nodes within the MCTS-Gen framework.
We conduct extensive experiments on the dataset constructed from TinyStories (Eldan and Li, 2023), featuring complex interaction with multiple characters across diverse backgrounds. Experimental results demonstrate AniMaker’s superior performance across VBench, AniEval, and human evaluation, with significantly improved multi-candidate generation efficiency. This validates AniMaker’s effectiveness in bringing AI-generated storytelling videos closer to production-grade quality (as shown in Figure 1).
In summary, our main contributions are in three key aspects:
  • We propose AniMaker, a fully automated multi-agent framework for generating coherent, multi-character, multi-scene animation from textual stories. The framework features MCTS-Gen, a novel search-based generation strategy that efficiently balances exploration and resource usage.
  • We develop AniEval, the first comprehensive evaluation framework specifically designed for multi-shot storytelling animation. It provides context-aware assessment by analyzing clips in relation to their surrounding content.
  • Experimental results demonstrate AniMaker’s superior performance across all evaluation frameworks—achieving best scores across all scene image metrics, ranking 1st in VBench, 14.6% higher scores in our AniEval, and better human ratings (3.22 versus 2.07).
Figure 2. The overall architecture of our AniMaker framework. Given a story input, Director Agent creates detailed scripts and storyboards with reference images. Photography Agent generates candidate video clips using MCTS-Gen, which optimizes exploration-exploitation balance. Reviewer Agent evaluates clips with our AniEval assessment system. Post-production Agent assembles selected clips, adds voiceovers, and synchronizes audio with subtitles. This multi-agent system enables fully automated, high-quality animated storytelling.
2. Related Work
Storyboard Visualization
Storyboard visualization bridges script to video by generating coherent image sequences, demanding flexible, reusable character-background modules for multiple scenes. While adapter-based techniques like IP-Adapter (Ye et al., 2023), Mix-of-Show (Gu et al., 2023), T2I-Adapter (Mou et al., 2024), ROICtrl (Gu et al., 2024), and StoryAdapter (Mao et al., 2024) realize character consistency, they often overlook background continuity. Consistency-aware methods try to address this challenge. StoryGen (Liu et al., 2024b) iteratively generates images using previous visual-language context, while StoryDiffusion (Zhou et al., 2024) employs a training-free Consistent Self-Attention module to improve feature alignment. Despite these advancements, the precise preservation of visual details, especially with multiple characters or changing backgrounds, remains a challenge.
Video Generation
Driven by the success of diffusion models, VDM (Ho et al., 2022) pioneers their application to video generation, leading to improved models like Stable Video Diffusion (Blattmann et al., 2023), ModelScope (Wang et al., 2023), VideoCrafter1 (Chen et al., 2023), and VideoCrafter2 (Chen et al., 2024). While these models excel at animating single text or image, they struggle with maintaining consistency across multiple video clips. Following Sora(Liu et al., 2024c), research has been increasingly focusing on coherent storytelling videos, with several works(Wu et al., 2024; Kim et al., 2024; Dalal et al., 2025; Guo et al., 2025) enhancing temporal modeling and contextual understanding. Despite these advancements, challenges persist as single-pass long video generation often suffers from inconsistencies and repetitive content over time. Considering this, models like HunyuanVideo (Kong et al., 2024), Wan (Wan et al., 2025), and CogVideoX (Yang et al., 2024) still prioritize the effective utilization of text prompts and reference images, resulting in more faithful generation and improved motion depiction. Industry solutions (Bao et al., 2024; Pika Labs, 2025; Pixverse AI, 2025; Runway AI, Inc., 2025) extend these advancements, enabling features like start/end frame-specific control.
Storytelling Video Agent
Storytelling video generation typically follows a pipeline: script → scene images → video clips → final video composition. Large multimodal models (LMMs) often act as high-level planners, coordinating each stage. Models such as VideoDirectorGPT (Lin et al., 2023), DreamStory (He et al., 2024), AnimDirector (Li et al., 2024), MM-StoryAgent (Xu et al., 2025), MovieAgent (Wu et al., 2025), and DreamFactory (Xie et al., 2024) adhere to this conventional pipeline. Notably, certain tasks like scriptwriting, character design, voiceovers, and sound effects are typically handled manually during pre- or post-processing.
3. Methodology
Our proposed framework, AniMaker, automates the creation of storytelling animation from text. It uses a multi-agent system that mirrors professional animation pipelines, incorporating novel components for efficient candidate generation and comprehensive evaluation. The overall architecture is detailed in Figure 2.
3.1. Task Formulation
Automated generation of storytelling animation $\mathbf{V}_{\text{final}}$, from text input $\mathbf{T}_{\text{prompt}}$, can be represented as:
$\mathcal{F}:\mathbf{T}_{\text{prompt}}\rightarrow\mathbf{V}_{\text{final}}$
This transformation is realized by mapping through several crucial intermediate representations:
  • Script ($\boldsymbol{\mathcal{P}}_{\text{script}}$): Derived from $\mathbf{T}_{\text{prompt}}$, the script
    $\boldsymbol{\mathcal{P}}_{\text{script}}=((\mathbf{shot}_{k})_{k=1}^{\mathbf{N}_{\text{clips}}},\mathbf{K}_{cut})$
    defines:
    • An ordered sequence of $N_{\text{clips}}$ shots, where each $\mathbf{shot}_{k}=(\mathbf{d}_{k},\mathbf{C}_{k},\mathbf{B}_{k})$ specifies its textual description $\mathbf{d}_{k}$, involved characters $\mathbf{C}_{k}$, and background $\mathbf{B}_{k}$.
    • A set of indices $\mathbf{K}_{cut}\subseteq\{1,\dots,\mathbf{N}_{\text{clips}}\}$, marking shot transitions, indicating where KeyFrame guidance is required for the new shot.
  • Storyboard($\boldsymbol{\mathcal{S}}_{\text{board}}$): Storyboard is established from $\boldsymbol{\mathcal{P}}_{\text{script}}$, containing:
[…]
    • KeyFrames: for each $\mathbf{shot}_{k}$ where $k\in\mathbf{K}_{cut}$, a KeyFrame $\mathbf{F}_{k}^{\text{key}}$ is defined:
      $\mathbf{F}_{k}^{\text{key}}=\text{Visualize}(\{\mathbf{I}_{c}^{\text{char}}\mid\boldsymbol{c}\in\mathbf{C}_{k}\},\mathbf{I}_{\mathbf{B}_{k}}^{\text{bg}},\mathbf{d}_{k})$
      where $\text{Visualize}(\cdot)$ generates a keyframe by integrating the specified characters, background, and text description into a multimodal prompt.
  • Video Clip Sequence ($\mathbf{v}$): $\mathbf{v}=(\mathbf{v}_{1},\dots,\mathbf{v}_{N_{\text{clips}}})$ is generated, corresponding to $(\mathbf{shot}_{k})_{k=1}^{N_{\text{clips}}}$. Let $G_{K}$ and $G_{C}$ be abstract generative processes of each clip $\mathbf{v}_{k}$:
    • If $k\in\mathbf{K}_{cut}$ (after shot transition): $\mathbf{v}_{k}\sim G_{K}(\mathbf{F}_{k}^{\text{key}},\mathbf{d}_{k})$.
    • If $k\notin\mathbf{K}_{cut}$: $\mathbf{v}_{k}\sim G_{C}(\text{last\_frame}(\mathbf{v}_{k-1}),\mathbf{d}_{k})$.
  • Final Video ($\mathbf{V}_{\text{final}}$): The assembled clip sequence $\text{Assemble}(\mathbf{v})$ then undergoes post-processing, utilizing information from $\boldsymbol{\mathcal{P}}_{\text{script}}$ to produce the polished $\mathbf{V}_{\text{final}}$. This may include additions like voiceovers and subtitles.
    $\mathbf{V}_{\text{final}}=\text{PostProcess}(\text{Assemble}(\mathbf{v}),\boldsymbol{\mathcal{P}}_{\text{script}})$
3.2. Pipeline Overview
AniMaker transforms textual input ($\mathbf{T}_{\text{prompt}}$) into compelling storytelling animation ($\mathbf{V}_{\text{final}}$) through four specialized agents working collaboratively. The Director Agent (Section 3.3.1) creates a detailed storyboard ($\boldsymbol{\mathcal{S}}_{\text{board}}$), the Photography Agent (Section 3.3.2) generates candidate video clips ($\mathbf{v}_{k}$) using MCTS-Gen (efficiently exploring generation space conserving computational resources), the Reviewer Agent (Section 3.3.3) evaluates clips with AniEval (our context-aware evaluation framework), and the Post-production Agent (Section 3.3.4) assembles clips with voiceover.
3.3. Multi-Agent Framework: AniMaker
3.3.1. Director Agent
[…]
3.3.2. Photography Agent
Converting storyboards to clips in multi-shot AI video generation presents challenges including distorted appearances, inconsistent motion, and object inconsistencies. Inspired by filmmaking’s ”NG” concept and ”gacha” methods prevalent in AIGC, we recognize the necessity of generating multiple candidate clips to select the optimal one. For optimal selection, each clip must not only possess high individual quality but also ensure consistency and coherence with both preceding and succeeding clips. However, naively generating $k$ candidates per clip creates a combinatorial explosion (e.g., $k^{2}$ for two-clip sequences). Notably, poor-quality current clips allow for pruning the search space, as further exploration down such paths is unlikely to be satisfactory.
Thus we propose MCTS-Gen, a Monte Carlo Tree Search (MCTS)-inspired method for multi-clip video generation (Figure 3). MCTS naturally fits this task: multi-clip sequences correspond to tree paths, with each clip as a node. Crucially, clip evaluation considers both intrinsic quality and inter-clip consistency, aligning with MCTS’s backpropagation where child’s scores update the parent’s evaluation results. MCTS-Gen iteratively constructs a chosen path of selected video clips, extending this path by one clip per iteration. This process is controlled by the following parameters: $\mathbf{w_{1}}$ (initial candidate count per node), $\mathbf{w_{2}}$ (UCT-guided expansion times per iteration), and $\boldsymbol{\alpha}$ (exploration-exploitation balancing factor, default 1). Specifically, we use Wan 2.1 (Wan et al., 2025) for video clip generation, and the algorithm proceeds as follows:
1. Expansion: $\mathbf{w_{1}}$ (3 in Figure 3) initial child clips are generated from the chosen path’s terminal node (node 1 in Figure 3). These clips (node 3, 4, 5 in Figure 3) are then scored using AniEval (detailed in Section 3.3.3) and ranked.
2. Simulation: Further $\mathbf{w_{2}}$ (3 in Figure 3) expansions are generated guided by the UCT score:
$\boldsymbol{UCT}(\boldsymbol{node_{j}})=\frac{2.0}{\boldsymbol{rank}(\boldsymbol{node_{j}})+1}+\boldsymbol{\alpha}\cdot\sqrt{\frac{2.0}{\boldsymbol{child\_count}(\boldsymbol{node_{j}})+1}}$
where $\boldsymbol{rank}$ is from the initial AniEval scoring, $\boldsymbol{child\_count}$ is dynamically updated, and $\boldsymbol{\alpha}$ balances exploitation and exploration. The node with the highest UCT score generates a new child clip.
3. Backpropagation: AniEval scores of child clips (nodes 6, 7, 8 in Figure 3) propagate upwards. A parent node’s score is updated by adding the average score of children to its own.
4. Selection: The node with the highest AniEval score (node 3 in Figure 3) is added to the chosen path, then generating new child clips until reaching a total of $\mathbf{w_{1}}$ children, allowing the iterative generation process to continue.
This MCTS-Gen process systematically balances the exploration of diverse video generation space with the exploitation of promising paths, aiming to efficiently construct a high-quality, coherent video sequence while managing computational resources.
Figure 3. Illustration of our MCTS-Gen strategy for efficient Best-of-N Sampling.
3.3.3. Reviewer Agent
Existing objective metrics, such as CLIP Score and Inception Score, may identify superior video generation models, but often struggle to differentiate among candidates from the same model using the same prompt. Similarly, the widely adopted VBench has significant limitations. For instance, some of its metrics, like ”dynamic degree,” are overly simplistic, merely measuring pixel changes rather than accurately reflecting character action. More critically, VBench’s ”Consistency” metrics, based on single-clip segmentation, prove unsuitable for multi-shot animation, which inherently involves frequent character and scene changes.
To address evaluation challenges, we introduce AniEval, a comprehensive evaluation framework built on EvalCrafter (Liu et al., 2024a). AniEval refines EvalCrafter’s metrics for fully automated evaluation, e.g., by automating action assessment through comparing prompted character actions with those identified in video clips.
Furthermore, responding to the specific demands of evaluating multi-shot animation characterized by multiple characters and diverse scenes, AniEval introduces several additional metrics: DreamSim (Fu et al., 2023) assesses overall frame consistency; Count-Score (Cheng et al., 2024) aims at the issue of objects appearing or disappearing between shots; Face Consistency evaluates animated character facial consistency by training an InceptionNext (Yu et al., 2024) model on the Anime Face Dataset (splcher, 2019), overcoming the limitations of conventional face recognition methods like MTCNN (Ku and Dong, 2020) in anime face detection and tracking.
In conclusion, AniEval comprises 4 primary domains with 14 fine-grained metrics for comprehensive assessment (Table 1). Additionally, AniEval supports contextual scoring by evaluating clips based on preceding and succeeding content, providing robust evaluation for multi-shot animation generation.
Table 1. AniEval Metrics.
Domain Metric Brief Description
Overall Video Quality VQA_A Aesthetic video quality
VQA_T Technical video quality
MusIQ Frame quality score
Text-Video Alignment Text-Video Consistency Measured by CLIP
Text-Story Consistency Measured by BLIP-BLEU
Detection-Score Object generation accuracy
Count-Score Key object count accuracy
Video Consistency DreamSim Perceptual frame-frame similarity
Face Consistency Animated character
facial consistency
Warping Error Temporal inconsistency
via pixel differences
Semantic Consistency Temporal semantic coherence
via CLIP
Motion Quality Action Recognition Actions-prompt consistency
Action Strength Motion intensity via Flow-Score
Motion AC-Score Motion amplitude-prompt
consistency
3.3.4. Post-production Agent
The Post-production Agent transforms video clip sequences into a polished animation film through three stages. Firstly, it leverages Gemini 2.0 Flash to generate a detailed voiceover script specifying narration, dialogue, emotional tones, and desired voice timbres. The agent then selects appropriate voice profiles based on character attributes (age, gender) and assesses text length for audio-visual synchronization. Secondly, the script is processed through CosyVoice2 (Du et al., 2024) to generate audio tracks, which undergo verification for duration consistency and content accuracy. Finally, the agent employs the MoviePy library for film assembly, integrating validated subtitles and performing comprehensive editing to ensure precise synchronization between visuals, voiceover, and subtitles.
4. Experiments
4.1. Settings
4.1.1. Datasets
To evaluate our AniMaker, we sample 10 narratives from TinyStories (Eldan and Li, 2023). These narratives feature complex multi-character interactions across diverse backgrounds, providing an ideal testbed for multi-shot animation generation.
4.1.2. Baseline
We evaluate several state-of-the-art storytelling models: StoryGen, StoryDiffusion, and StoryAdapter (visual narrative specialists), alongside MovieAgent, MMStoryAgent, and VideoGen-of-Thought (video generators). The latter group utilizes their built-in video modules, while StoryDiffusion and StoryAdapter are paired with external image-to-video models (CogVideoX and Wan 2.1).
Table 2. Scene Image evaluation on Contextual Coherence (Coherence), Image-Image Similarity (I-I Sim), and Text-Image Similarity (T-I Sim).
[…]
StoryGen 0.54 0.77 0.22
StoryDiffusion 0.70 0.80 0.25
StoryAdapter 0.78 0.83 0.25
MovieAgent 0.59 0.65 0.23
MMStoryAgent 0.78 0.83 0.26
VideoGen-of-Thought 0.71 0.77 0.23
AniMaker(Ours) 0.81 0.83 0.31
4.1.3. Evaluation Metrics
For scene image generation, we evaluate text-to-image alignment and cross-image consistency. Metrics include Text-to-Image CLIP (Coherence) (Radford et al., 2021), Image-to-Image Similarity (I-I Sim) (Gal et al., 2022), and Text-Image Similarity (T-I Sim) (Hessel et al., 2021). Video generation is assessed using VBench and our AniEval for comprehensive evaluation.
4.2. Qualitative Analysis
Based on samples in Figures 5 and 6, we qualitatively analyze AniMaker’s output, focusing on visual fidelity and narrative coherence in multi-character, multi-scene animated storytellings.
Enhanced Consistency
AniMaker achieves superior visual consistency across scenes. While baselines struggle with character and background consistency during transitions, AniMaker maintains visual characteristics even when switching between different scenes. This stems from the Director Agent’s storyboard creation, providing consistent reference images for characters and backgrounds. Additionally, AniEval ensures coherent clip selection by evaluating cross-clip consistency metrics.
Improved Action Representation
AniMaker excels at depicting complex and extended character actions. Baseline methods often produce incomplete movements, particularly for multi-step action sequences (e.g., squatting, picking up an object, standing up, and walking away in Figure 5). Our MCTS-Gen strategy enables the Photography Agent to explore and select clip sequences that concatenate into coherent, complete long actions.
Seamless Transitions Between Clips
AniMaker delivers smoother video transitions through effective generation and selection mechanisms. The Reviewer Agent uses AniEval, integrating cross-clip consistency metrics like DreamSim and applying contextual scoring across adjacent clips, minimizing visual disruptions.
In summary, AniMaker outperforms existing methods through superior visual consistency across scenes, effective depiction of complex action sequences, seamless inter-clip transitions, and robust handling of multi-character, multi-scene narratives.
Table 3. VBench evaluation results, presenting scores for Image Quality (I.Q.), Semantic Consistency (S.C.), Background Consistency (B.C.), Animation Quality (A.Q.), Motion Smoothness (M.S.), Dynamic Degree (D.D.), and Average Rank (Rk. Avg. - the average ranking position across all models).
Model I.Q.$\uparrow$ S.C.$\uparrow$ B.C.$\uparrow$ A.Q.$\uparrow$ M.S.$\uparrow$ D.D.$\uparrow$ Rk. Avg.$\downarrow$
StableDiffusion+Cog. 75.52 78.05 85.27 59.61 97.63 33.58 5.83
StableDiffusion+Wan. 76.93 78.54 87.43 69.75 96.67 60.30 4.33
StoryAdapter+Cog. 76.17 72.17 88.03 63.55 98.16 26.71 5.33
StoryAdapter+Wan. 75.96 75.04 88.64 73.38 97.02 84.73 4.17
MovieAgent 72.09 68.61 79.84 55.40 99.01 35.44 6.33
MMStoryAgent 76.41 87.27 90.74 73.84 99.80 0.00 2.67
VoT 63.85 75.11 85.78 74.91 99.25 3.50 4.83
AniMaker(Ours) 76.96 84.27 89.06 69.79 98.50 66.97 2.50
4.3. Quantitative Comparisons
Scene Image Generation Analysis
Scene images (Keyframes) are the core output of a storyboard. Table 2 demonstrates that AniMaker outperforms all competing methods across metrics. Notably, AniMaker achieves a Text-to-Image Similarity (T-I Sim) score of 0.31, representing a 19.2% improvement over the best-performing baseline. This advantage stems from our multimodal storyboard generation approach that incorporates character references, background references, and text prompts for keyframe generation, rather than relying solely on text.
VBench Evaluation Analysis
Table 3 shows that AniMaker achieves the best average rank (2.50), demonstrating consistent top-tier performance across metrics. While MMStoryAgent excels in Semantic Consistency (S.C.) and Background Consistency (B.C.), its 0.00 score in Dynamic Degree (D.D.) reveals a critical limitation—it produces static image sequences resembling comic strips rather than true animation. This exposes VBench’s limitations: it favors static consistency (where even brief static images can achieve high scores) and focuses on individual clips rather than multi-scene, multi-character animation. These findings highlight the need for a new evaluation framework specifically designed for storytelling animation quality.
AniEval Evaluation Analysis
We introduce AniEval for better evaluation of storytelling Animation. Table 4 presents results from our AniEval framework. AniMaker outperforms all competitors with a total score of 76.72, representing a 14.6% improvement over the second-best model (VideoGen-of-Thought at 66.93). Particularly noteworthy is AniMaker’s exceptional performance in Video Consistency (V.C.), surpassing the best-performing baseline by 15.5%. AniMaker’s relatively lower performance in Text-Video Alignment (T.V.A.) can be attributed to our agent framework’s creative adaptation of stories into scripts, where additional narrative elements are introduced. Compared to VBench, AniEval’s assessment results align remarkably better with human evaluations (Table 5), confirming that AniEval provides a more comprehensive and accurate assessment of multi-shot animation quality than previous metrics.
MCTS-Gen Parameter Analysis
Figure 4 shows how MCTS-Gen parameters $\mathbf{w_{1}}$ (initial candidate count) and $\mathbf{w_{2}}$ (expansion iterations) affect generation quality. Two patterns emerge: higher $\mathbf{w_{1}}$ values improve AniEval scores by providing more initial candidates, while higher $\mathbf{w_{2}}$ values enhance performance by better evaluating each clip’s suitability for continuous generation. Importantly, once certain thresholds are reached, configurations with fewer total generations (e.g., $\mathbf{w_{1}}=3,\mathbf{w_{2}}=3$ with 4.37 generations per node) perform comparably to those requiring more (e.g., $\mathbf{w_{1}}=3,\mathbf{w_{2}}=5$ with 5.76 generations per node). This demonstrates MCTS-Gen’s efficiency—compressing the search space by over 50% compared to exhaustive search strategies (which require 9 generations per node for the same candidates) while maintaining quality.
Table 4. AniEval evaluation results, presenting scores for Overall Video Quality (O.V.Q.), Text-Video Alignment (T.V.A.), Video Consistency (V.C.), Motion Quality (M.Q.), and Total Performance.
[…]
StoryDiffusion+CogVideoX 46.54 86.05 47.14 70.35 56.75
StoryDiffusion+Wan 2.1 47.07 84.99 47.13 71.00 56.55
StoryAdapter+CogVideoX 56.76 87.38 55.89 69.73 63.95
StoryAdapter+Wan 2.1 60.39 86.99 51.41 72.11 62.37
MovieAgent 41.17 68.50 68.68 70.16 61.95
MMStoryAgent 47.93 75.27 63.54 61.39 62.79
VideoGen-of-Thought 66.17 72.95 65.42 66.72 66.93
AniMaker(Ours) 81.87 74.30 79.35 72.66 76.72
Figure 4. AniEval Score of Different $\mathbf{w_{1}}$ (initial candidate count) and $\mathbf{w_{2}}$ (expansion iterations) Combinations.
4.4. Human Rating
Following MovieAgent’s evaluation framework (Wu et al., 2025), we conduct human evaluation with 10 participants on 90 storytelling videos from 9 models across 10 stories. Each video was rated on a 1-5 scale across five dimensions: Visual Appeal, Script Faithfulness, Narrative Coherence, Character Consistency, and Physical Law Adherence. Note that evaluating complete storytelling videos (rather than individual clips) typically yields lower scores due to the increased complexity and length. As shown in Table 5, our model achieves superior performance across all metrics, with particularly strong results in Character Consistency.
Table 5. Human rating results, presenting scores for Character Consistency (C.C.), Narrative Coherence (N.C.), Physical Law Adherence (P.L.), Script Faithfulness (S.F.), and Visual Appeal (V.A.) using a 1-5 scale.
Model C.C.$\uparrow$ N.C.$\uparrow$ P.L.$\uparrow$ S.F.$\uparrow$ V.A.$\uparrow$ Avg.$\uparrow$
StoryDiffusion+CogVideoX 1.37 1.48 1.37 1.67 1.56 1.49
StoryDiffusion+Wan 2.1 2.00 1.82 1.82 2.00 2.11 1.95
StoryAdapter+CogVideoX 1.64 1.39 1.46 1.68 1.71 1.58
StoryAdapter+Wan 2.1 2.04 1.82 1.89 2.04 2.57 2.07
MovieAgent 1.19 1.26 1.44 1.26 1.48 1.33
MM_StoryAgent 1.62 1.62 1.72 1.83 2.24 1.81
VideoGenoT 1.67 1.74 1.78 1.74 2.26 1.84
AniMaker(Ours) 3.44 3.24 3.04 3.08 3.28 3.22
4.5. Ablation Studies
We conduct two key ablation experiments. First, we ablate MCTS-Gen by setting $w_{1}=1,w_{2}=1$, essentially generating only one candidate per clip. This change results in a 7.1% reduction in performance (the green triangle in Figure 4) on AniEval, confirming the importance of our MCTS-driven generation strategy. Notably, even this ablated version still outperforms the best-performing baseline method by 6.6%, further demonstrating that our multi-agent framework remains highly competitive even without clip candidate selection. Next, we ablate AniEval by generating five candidates per clip with VBench for selection. This yields a score of 73.18, a 4.6% decrease compared to our full method. Qualitative assessment of the resulting videos also reveals noticeable degradation in action expressiveness and cross-clip consistency, highlighting the importance of AniEval for storytelling animation.
5. Limitations
Despite employing state-of-the-art models for storyboard and video clip generation, a gap remains between current model capabilities and commercial film production quality. Key limitations include poor adherence to physical laws and unrealistic interactions between scene elements, creating bottlenecks in production-ready output. However, our framework adopts a modular, plug-and-play architecture that enables seamless model integration. As more advanced models emerge, we will continuously update our framework and share these workflow improvements with the community, especially mitigating the hallucination of models (Wang et al., 2024).
6. Conclusion
We present AniMaker, a comprehensive multi-agent framework that transforms text input into coherent storytelling animation by emulating professional workflows. Our system introduces two key innovations: MCTS-Gen, which optimizes exploration-exploitation balance during clip generation, and AniEval, the first evaluation framework specifically designed for multi-shot storytelling animation. AniMaker orchestrates specialized agents that seamlessly collaborate across storyboarding, generation, evaluation, and post-production stages. Our quantitative results validate the effectiveness of this approach, with substantial gains in both technical metrics and perceived quality. These advances mark an important step toward bridging the gap between AI-generated content and professional animation standards, paving the way for more accessible and high-quality animated storytelling production.